# The Honest Truth About Causal Trees: Accuracy Limits for Heterogeneous Treatment Effect Estimation

### v1

@@ 4 Main Results
The following theorem summarizes our first main result. Let $e$ denote Euler’s constant.
@@
Suppose Assumption 1 holds, and the underlying causal tree has at least one split (i.e., at least two terminal nodes). Then, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and all $b\in(0,1)$,
@@
Section 4.1 gives an overview of the proof strategy of Theorem 1, with all omitted technical details given in the supplemental appendix (see Section SA-1.2 for details). Our proof relies on several non-asymptotic approximation steps for the suprema of partial sums and various Gaussian processes leveraging key technical results from Chernozhukov et al. (2017), Chernozhuokov et al. (2022), Csörgö and Révész (1981), Csörgö and Horváth (1997), Eicker (1979), El-Yaniv and Pechyony (2009), Göing-Jaeschke and Yor (2003), Horváth (1993), Latała and Matlak (2017), Petrov (2007), Shorack and Smythe (1976), and Skorski (2023). As a technical by-product, we correct a mistake in Eicker (1979): see Remark SA-1 in the supplemental appendix.
Theorem 1 presents precise lower bounds on the uniform convergence rate of the six causal tree estimators introduced in Section 2. Starting with procedures that do not employ sample splitting, Theorem 1 demonstrates that the three estimators $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{DIM}}(\mathbf{x})$, $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{IPW}}(\mathbf{x})$ and $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{SSE}}(\mathbf{x})$ cannot achieve a uniform convergence rate of $n^{-b/2}\sqrt{\log\log n}$, for any $b>0$. That is, they must have a worse than polynomial-in-$n$ uniform convergence rate, and thus suffer from low accuracy in estimating heterogeneous treatment effects in certain regions of the support $\mathcal{X}$.
Athey and Imbens (2016), and many others, argue that sample splitting (the so-called “honesty” property) can improve the performance of machine learning estimators, and in particular their proposed causal tree estimators, because such sample usage de-couples the causal tree construction and the CATE estimation steps. The second result in Theorem 1 considers exactly their honest causal tree estimators, $\hat{\tau}^{\mathtt{HON}}_{\mathtt{DIM}}(\mathbf{x})$, $\hat{\tau}^{\mathtt{HON}}_{\mathtt{IPW}}(\mathbf{x})$ and $\hat{\tau}^{\mathtt{HON}}_{\mathtt{SSE}}(\mathbf{x})$. It follows from the theorem that these estimators cannot achieve a uniform convergence rate that is polynomial-in-$n$ either. Notably, our results show that sample splitting (or honesty) improves the best achievable uniform convergence rate of the estimators, but this improvement is quite modest: the penalty term $\sqrt{\log\log n}$ is removed, thereby improving the uniform convergence rate by a very slow factor.
The results in Theorem 1 offer a pessimistic outlook on the utility of adaptive decision tree methods in causal inference when the goal is to learn about heterogeneous treatment effects: the estimators cannot perform well pointwise (and hence uniformly) over the entire support of the covariates; see Section 4.1 for more formal details. As a point of contrast, the same procedures considered in Theorem 1 can achieve near-optimal convergence rates “on average” over $\mathcal{X}$, as the following theorem establishes. Here again, honesty delivers only negligible improvements of order $\log(p)$.
@@
Suppose Assumption 1 holds and the underlying causal tree has depth at most $K\geq 1$, and let $F_{\mathbf{X}}(\mathbf{x})=\mathbb{P}(\mathbf{x}_{i}\leq\mathbf{x})$. Then, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$,
@@
The proof of this theorem is given in the supplemental appendix (see Section SA-1.2 for details). It leverages ideas and technical results from Györfi et al. (2002) and Klusowski and Tian (2024). Crucially, the result applies only when Assumption 1 holds, that is, when $\tau(\mathbf{x})$ is constant. The main purpose of Theorem 2 is to demonstrate that in the same basic setting when uniform convergence fails, causal decision trees nonetheless achieve favorable performance on average in an integrated mean‐squared sense. A natural way to interpret the juxtaposition between Theorem 1 and Theorem 2 is related to the often claimed tension between causal inference and prediction in machine learning settings: adaptive causal trees can perform poorly pointwise (hence uniformly), but excellently on average, over the feature space.
From a technical perspective, the results in Theorem 2 are new in the context of causal tree estimation and, notably, for the formal comparison between no-sample-splitting and honest implementations. Furthermore, our theoretical work in the supplemental appendix establishes the integrated mean‐squared error bounds with high-probability, enabling a sharper comparison with Theorem 1. For example, for the case of no-sample-splitting, we show that
@@
where $C_{1}$ is the constant in Theorem 2.
@@ 4.1 Proof Strategy of Theorem 1
Underlying our theoretical insights are a collection of technical results concerning a decision stump, and hence a decision tree of depth one. For each tree splitting criteria and sample splitting design, we first study the probabilistic properties of the split location at the root node, and thus characterize the regions of the support $\mathcal{X}$ where the first split index is most likely to realize. These theoretical results also characterize the effective sample size of the resulting child nodes. We establish that with non-vanishing probability, the first split will concentrate near a region of the boundary of the parent node (a cell in the partition of $\mathcal{X}$), from the beginning of any tree construction. More precisely, let $\hat{\imath}=n(\mathsf{t}_{L})$ and $\hat{\jmath}$ be the CART split index and split variable at the root node, respectively, with $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$, noticing that the first split coincide for no-sample-splitting and honest constructions. For each $a,b\in(0,1)$ with $a<b$ and $j\in\{1,2,\dots,p\}$, and $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$, we have
$\liminf_{n\to\infty}\mathbb{P}\big{(}n^{a}\leq\hat{\imath}\leq n^{b},\;\;\hat{\jmath}=j\big{)}=\liminf_{n\to\infty}\mathbb{P}\big{(}n-n^{b}\leq\hat{\imath}\leq n-n^{a},\;\;\hat{\jmath}=j\big{)}\geq\frac{b-a}{2pe}.$ (6)
@@
The slow uniform convergence rate of a decision stump estimator occurs because the optimal split point concentrates near the boundary of the support, causing the two nodes in the stump to be imbalanced, with one containing a much smaller number of samples, and therefore rendering a situation where local averaging is less accurate. This can be deduced from (6): for each coordinate $j=1,2,\dots,p$ and $b\in(0,1)$, there is non-vanishing $b/(pe)$ probability that the child cells $\{\mathbf{x}\in\mathcal{X}:x_{j}\leq\hat{\varsigma}\}$ or $\{\mathbf{x}\in\mathcal{X}:x_{j}>\hat{\varsigma}\}$ are highly anisotropic and will contain at most $n^{b}$ samples. Thus, with non-vanishing probability, the causal tree procedures will exhibit arbitrarily slow convergence rate in a region of $\mathcal{X}$. These results are then carefully recycled to characterize the properties of the deeper trees: due to their recursive nature, and since $p>1$, the problematic regions take the form of many hyper-rectangles, and will realize anywhere in $\mathcal{X}$, with non-vanishing probability.
@@
The core of proof strategy is to study the tree construction as the maximizer of the split criterion from (4) and (5), as indexed by the optimal split location and covariate coordinate. We leverage non-asymptotic high-dimensional central limit theorems, Gaussian comparison inequalities, Gaussian process embeddings, the Darling-Erdös theorem, and empirical process techniques (El-Yaniv and Pechyony, 2009; Petrov, 2007; Shorack and Smythe, 1976; Skorski, 2023), as explained in the following four main steps.
@@
Step 1: Split Criterion Approximation. Using empirical process theory techniques, we establish an asymptotic equivalence between the split criterion underlying each of the causal tree estimators and the split criterion of a standard (non-causal) decision regression tree employing CART. For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, the latter can be viewed as a standard regression tree with transformed outcomes $y_{i}\frac{d_{i}-\xi}{\xi(1-\xi)}$. For $l=\mathtt{SSE}$, approximating process is the sum of two independent split criterion processes, one with transformed outcome $\frac{d_{i}}{\xi}y_{i}$ for treated units, and the other with transformed outcome $\frac{1-d_{i}}{1-\xi}y_{i}$ for control units. We employ a careful truncation argument to remove extremely small or large split indices (Csörgö and Horváth, 1997, Theorem A.4.1), where empirical process techniques are hard to apply.
@@
Step 2: Conditional Gaussian Approximation. We show that, conditional on the covariates ordering, the square root of the split criterion processes from step 1 can be approximated by Gaussian processes with the same conditional covariance structure. For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, we view the split criterion process as a summation of i.i.d. high-dimensional random vectors, each entry corresponding to one pair of split index and coordinate. The high-dimensional central limit theorem of (Chernozhukov et al., 2017, Theorem 2.1) implies that the split criterion process in high-dimensional vector form is close to a high-dimensional Gaussian random vector with the same covariance matrix conditional the ordering, the latter can then be interpreted as a Gaussian process conditional on the ordering. Due to the structure of the splitting criteria, a high-dimensional CLT for hyper-rectangles is sufficient. For $l=\mathtt{SSE}$, we stack the control and treatment groups process in a twice as long high-dimensional vector. However, due to the structure the splitting criteria in this case, we employ instead Chernozhukov et al. (2017, Proposition 3.1), which gives a high-dimensional CLT for convex sets.
Step 3: Unconditional Gaussian Approximation. For the special case of $p=1$, this step is not necessary because there is only one ordering possible. However, for $p>1$, recursive decision trees find the best split along each dimension of $\mathbf{x}_{i}$, which implies a different ordering of the vector. Nevertheless, we show that the conditional Gaussian process from step 2 is close to an unconditional Gaussian process with zero correlation for different split coordinate indexes. Zero correlation between splits of different coordinates implies that the (sub)-processes corresponding to splitting different coordinates are asymptotically independent, reducing the problem to studying the $\operatorname*{arg\,max}$ of the split criterion over one coordinate. The result is proven by applying a Gaussian-to-Gaussian comparison inequality (Chernozhuokov et al., 2022, Proposition 2.1), after establishing an upper bound on the matrix $\max$ norm of the difference between the conditional covariance matrix (which depends on the ordering) and the unconditional covariance matrix (which does not depend on the ordering). For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, the results is immediate because the high-dimensional CLT was established over hyper-rectangles. For $l=\mathtt{SSE}$, the additional error induced by considering a simple convex sets approximation is be controlled using Nazarov’s inequality (Nazarov, 2003).
@@
Step 4: Lower bound on imbalanced split probability. The unconditional Gaussian approximation processes from Step 3 take the form of the square Euclidean norm of a univariate (for $l\in\{\mathtt{DIM},\mathtt{IPW}\}$) or bivariate (for $l=\mathtt{SSE}$) Ornstein-Uhlenbeck process, where the split and time of Ornstein-Uhlenbeck process satisfies a one-to-one transformation (Csörgö and Révész, 1981; Göing-Jaeschke and Yor, 2003). Since Darling-Erdös (Eicker, 1979; Horváth, 1993) allows for calculation of the maximum of norm of an O-U process within any time interval, we can find the lower bound on the probability of split occurs with a small or large index from (6) with the help of Gaussian correlation inequality (Latała and Matlak, 2017, Remark 3 (i)). In turn, this characterizes precisely the effective sample sizes of each child node.
The remaining of our proofs leverage the technical insights above, applying then recursively to understand deeper tree constructions and the concentration in probability properties of the resulting CATE estimates.
@@ 5 $\mathbf{X}$-Adaptivity and Inconsistency
The estimators considered in Theorem 1 either employ the full sample in their entire construction, or they rely on a two-sample independent split (honesty), where one subsample is use for training the tree, and the other is used for estimation of the conditional average treatment effects. As discussed in Devroye et al. (2013), and references therein, $\mathbf{X}$-adaptivity offers a middle ground between the two sample usage designs considered in Definition 2: the tree construction and the final estimation step share the same covariates but each step employs different outcomes variables, that is, the two subsamples are independence conditional on the covariates.
We leverage the idea of $\mathbf{X}$-Adaptivity, and study causal tree estimators where the outcome variable and treatment indicator are independent across all levels of the tree construction and the final CATE estimation step, but the same covariates are used throughout. This $\mathbf{X}$-adaptive data design is of theoretical interest because it offers a bridge between no-sample-splitting and honesty. The following definition formalizes the construction of the $\mathbf{X}$-adaptive causal tree estimators.
@@
    The dataset $\mathcal{D}$ is divided into $K+1$ datasets $(\mathcal{D}_{\mathsf{T}_{1}},\ldots,\mathcal{D}_{\mathsf{T}_{K}},\mathcal{D}_{\tau})$, with sample sizes given by $(n_{\mathsf{T}_{1}},\ldots,n_{\mathsf{T}_{K}},n_{\tau})$, respectively, and satisfying $n_{\mathsf{T}_{1}}=\cdots=n_{\mathsf{T}_{K}}=n_{\tau}$ (possibly after dropping $n\mod K$ data points at random). For each of the datasets $\mathcal{D}_{j}=\{(y_{i},d_{i},\mathbf{x}_{i}^{\top}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$, $j=1,\ldots,K$, replace $\{(y_{i},d_{i}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$ with independent copies $\{(\tilde{y}_{i},\tilde{d}_{i}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$, while keeping the same $\{\mathbf{x}_{i}:i=1,\ldots,n_{\mathsf{T}_{j}}\}$.
@@
    The maximal decision tree of depth $K$, $\mathsf{T}^{l}_{K}(\mathcal{D}_{\mathsf{T}_{1}},\cdots,\mathcal{D}_{\mathsf{T}_{K}})$, is obtained by iterating $K$ times the $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ splitting procedures in Definition 2, each time splitting all terminal nodes until (i) the node contains a single data point $(y_{i},d_{i},\mathbf{x}_{i}^{\top})$, or (ii) the input values $\mathbf{x}_{i}$ and/or all $(d_{i},y_{i})$ within the node are the same.
@@
As in the previous cases, if the distributions of $\varepsilon_{i}(0)$ and $\varepsilon_{i}(1)$ are symmetric about zero, then the $\mathbf{X}$-adaptive estimators are unbiased: $\mathbb{E}[\hat{\tau}^{\mathtt{X}}_{l}(\mathbf{x};K)]=\tau$, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$.
@@
The theorem establishes uniform inconsistency of the $\mathbf{X}$-adaptive causal tree estimator so long as $K_{n}\gtrsim\log\log n$. To put this side rate restriction in perspective, if $n/K_{n}\approx 1\text{ billion}$ then $\log\log(10^{9})\approx 3$. Therefore, the inconsistency of the estimator will manifest as soon as $K_{n}\approx 3$, a shallow tree when compared to those commonly encountered in practice (even in settings with much more moderate sample sizes, that is, with $n$ much smaller than $K_{n}$ billions). This result also shows that the integrated mean square error (IMSE) of a uniformly inconsistent $\mathcal{X}$-adaptive causal tree estimator can nonetheless decay at the optimal $\sqrt{n}$ rate, up to a poly-logarithmic-$n$ factor. As demonstrated before, the performance of the causal tree estimators can vary widely depending on whether the input $\mathbf{x}$ is average or worst case.

### v2

@@ 4 Main Results
The following theorem establishes our main lower bound on the uniform accuracy of causal tree estimators. Let $e$ denote the base of the natural logarithm.
@@
Suppose Assumption 1 holds and the underlying causal tree has at least one split (i.e., at least two terminal nodes). Then, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ and all $b\in(0,1)$,
@@
Section 4.1 provides an overview of the proof strategy for Theorem 1, with all technical details deferred to the supplemental appendix (see Section SA-1.2). The proof relies on several non-asymptotic approximations for suprema of partial sums and Gaussian processes, building on results from Chernozhukov et al. (2017), Chernozhuokov et al. (2022), Csörgö and Révész (1981), Csörgö and Horváth (1997), Eicker (1979), El-Yaniv and Pechyony (2009), Göing-Jaeschke and Yor (2003), Horváth (1993), Latała and Matlak (2017), Petrov (2007), Shorack and Smythe (1976), and Skorski (2023). As a technical by-product, Appendix A corrects an error in Eicker (1979).
Theorem 1 establishes lower bounds on the uniform convergence rate of the six causal tree estimators introduced in Section 2. For procedures without sample splitting, the estimators $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{DIM}}(\mathbf{x})$, $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{IPW}}(\mathbf{x})$, and $\hat{\tau}^{\mathtt{NSS}}_{\mathtt{SSE}}(\mathbf{x})$ cannot achieve a uniform convergence rate of order $n^{-b/2}\sqrt{\log\log n}$ for any $b>0$. In particular, these estimators converge more slowly than any polynomial rate in $n$, implying that their accuracy must deteriorate in some regions of the covariate space $\mathcal{X}$.
Athey and Imbens (2016) and subsequent work argue that sample splitting (the so-called “honesty” property) can improve the performance of machine learning estimators by decoupling model selection from estimation. The second result in Theorem 1 analyzes the corresponding honest causal tree estimators, $\hat{\tau}^{\mathtt{HON}}_{\mathtt{DIM}}(\mathbf{x})$, $\hat{\tau}^{\mathtt{HON}}_{\mathtt{IPW}}(\mathbf{x})$, and $\hat{\tau}^{\mathtt{HON}}_{\mathtt{SSE}}(\mathbf{x})$. The theorem shows that these estimators also fail to achieve polynomial convergence rates. While honesty improves the attainable rate slightly by removing the $\sqrt{\log\log n}$ factor, the resulting improvement is not even logarithmic in sample size.
Theorem 1 therefore paints a pessimistic picture for adaptive decision tree methods when the goal is to estimate heterogeneous treatment effects uniformly over the covariate space. In contrast, the same estimators can achieve near-optimal estimation accuracy when their performance is measured on average over $\mathcal{X}$, as shown by the following result. Let $F_{\mathbf{X}}(\mathbf{x})=\mathbb{P}(\mathbf{x}_{i}\leq\mathbf{x})$.
@@
Suppose Assumption 1 holds and the underlying causal tree has depth at most $K\geq 1$. Then, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$,
@@
The proof of Theorem 2 is given in the supplemental appendix (Section SA-1.2) and builds on ideas from Györfi et al. (2002) and Klusowski and Tian (2024). Importantly, the result applies only under Assumption 1, that is, when the CATE function is constant. The purpose of this theorem is to highlight that, even in the same setting where uniform convergence fails, causal decision trees can still achieve favorable performance in an integrated mean-squared sense. It remains an open question whether near-optimal mean square convergence rates can be achieved over larger classes of functions by adaptive decision trees constructed using CART-type procedures.
A natural interpretation of the contrast between Theorem 1 and Theorem 2 relates to the often-discussed tension between causal inference and prediction in machine learning. Adaptive causal trees may perform poorly pointwise (and hence uniformly), yet perform well on average over the covariate space. This contrast highlights a fundamental limitation of adaptive recursive partitioning for learning heterogeneous treatment effects: while tree-based methods can deliver strong predictive performance on average, they may fail to provide reliable estimation guarantees uniformly over the covariate space.
From a technical perspective, the results in Theorem 2 are new in the context of causal tree estimation, particularly for the formal comparison between no-sample-splitting and honest implementations. Moreover, our theoretical analysis in the supplemental appendix establishes high-probability bounds for the integrated mean-squared error, allowing for a sharper comparison with Theorem 1 (see Corollaries SA-12, SA-14, SA-32, SA-34 and Theorems SA-22 and SA-24). For example, in the case of no sample splitting we show that
@@
for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$, where $C_{1}$ is the constant appearing in Theorem 2.
@@ 4.1 Proof Strategy of Theorem 1
The key idea behind the proof is that greedy recursive partitioning tends to select highly imbalanced splits with non-vanishing probability. In particular, even in the simplest case of a decision stump (a tree of depth one), the optimal split often occurs near the boundary of the covariate space, producing child nodes with very small sample sizes. These small cells lead to large estimation variance in some regions of $\mathcal{X}$, which ultimately prevents causal tree estimators from achieving polynomial uniform convergence rates.
Underlying our theoretical results are several technical properties of decision stumps, and hence trees of depth one. For each tree splitting criterion and sample-splitting design, we first analyze the probabilistic behavior of the split location at the root node. This analysis characterizes the regions of the covariate space $\mathcal{X}$ where the first split is most likely to occur and determines the effective sample sizes of the resulting child nodes.
Our results show that, with non-vanishing probability, the optimal split concentrates near the boundary of the parent node (a cell in the partition of $\mathcal{X}$). As a consequence, one of the child nodes may contain only a very small number of observations. This phenomenon arises at the very first step of the recursive tree construction and ultimately drives the slow uniform convergence rate. More precisely, let $\hat{\imath}=n(\mathsf{t}_{L})$ and $\hat{\jmath}$ be the CART split index and split variable at the root node, respectively, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$. Note that the first split coincide for no-sample-splitting and honest constructions. For each $a,b\in(0,1)$ with $a<b$ and $j\in\{1,2,\dots,p\}$, we establish that
$\liminf_{n\to\infty}\mathbb{P}\big(n^{a}\leq\hat{\imath}\leq n^{b}\mid\hat{\jmath}=j\big)=\liminf_{n\to\infty}\mathbb{P}\big(n-n^{b}\leq\hat{\imath}\leq n-n^{a}\mid\hat{\jmath}=j\big)\geq\frac{b-a}{2e}.$ (6)
@@
The slow uniform convergence rate of the decision stump estimator arises because the optimal split tends to concentrate near the boundary of the support, producing highly imbalanced partitions. In such cases, one child node contains only a small number of observations, making the corresponding local average estimator highly variable. Relation (6) quantifies this phenomenon: for each coordinate $j=1,\dots,p$ and each $b\in(0,1)$, since $\mathbb{P}(\hat{\jmath}=j)=1/p$ by symmetry, there is non-vanishing probability, $b/(pe)$, that one of the child cells $\{\mathbf{x}\in\mathcal{X}:x_{j}\leq\hat{\varsigma}\}$ or $\{\mathbf{x}\in\mathcal{X}:x_{j}>\hat{\varsigma}\}$ are highly anisotropic and will contain at most $n^{b}$ samples. Consequently, with non-vanishing probability the estimator exhibits arbitrarily slow convergence in some region of $\mathcal{X}$. These insights extend to deeper trees. Because the partition is constructed recursively and $p>1$, the problematic regions propagate throughout the tree and appear as many hyper-rectangles scattered across $\mathcal{X}$ with non-vanishing probability.
@@
The core of the proof studies the tree construction as the maximizer of the split criteria in (4) and (5), indexed by the split location and the covariate coordinate. The analysis relies on non-asymptotic high-dimensional central limit theorems, Gaussian comparison inequalities, Gaussian process embeddings, the Darling–Erdös theorem, and empirical process techniques (El-Yaniv and Pechyony, 2009; Petrov, 2007; Shorack and Smythe, 1976; Skorski, 2023). The argument proceeds in four main steps.
@@
Step 1: Split Criterion Approximation. Using empirical process techniques, we establish an asymptotic equivalence between the split criterion underlying each causal tree estimator and the split criterion of a standard regression tree employing CART. For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, this corresponds to a regression tree applied to the transformed outcomes $y_{i}\frac{d_{i}-\xi}{\xi(1-\xi)}$. For $l=\mathtt{SSE}$, the approximating process is the sum of two independent split criterion processes, one based on the transformed outcome $\frac{d_{i}}{\xi}y_{i}$ for treated units and one based on $\frac{1-d_{i}}{1-\xi}y_{i}$ for control units. A truncation argument removes extremely small or large split indices where empirical process approximations are less reliable (Csörgö and Horváth, 1997, Theorem A.4.1).
@@
Step 2: Conditional Gaussian Approximation. Conditional on the ordering of the covariates, the square root of the split criterion process can be approximated by a Gaussian process with the same conditional covariance structure. For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, the split criterion can be written as a sum of i.i.d. high-dimensional random vectors indexed by split location and coordinate. Applying the high-dimensional central limit theorem of (Chernozhukov et al., 2017, Theorem 2.1), we obtain a Gaussian approximation conditional on the ordering. Because of the structure of the split criterion, a high-dimensional central limit theory over hyper-rectangles suffices. For $l=\mathtt{SSE}$, the treated and control components are stacked into a higher-dimensional vector and a central limit theory for convex sets (Chernozhukov et al., 2017, Proposition 3.1) is employed.
Step 3: Unconditional Gaussian Approximation. When $p>1$, different covariate coordinates induce different orderings of the observations. We therefore show that the conditional Gaussian process from Step 2 is close to an unconditional Gaussian process in which splits across different coordinates are asymptotically uncorrelated. This implies asymptotic independence of the corresponding subprocesses and reduces the problem to studying the maximization of the split criterion along a single coordinate. The approximation is established using a Gaussian comparison inequality (Chernozhuokov et al., 2022, Proposition 2.1) together with bounds on the difference between the conditional and unconditional covariance matrices. For $l=\mathtt{DIM}$ and $l=\mathtt{IPW}$, the argument follows directly from a high-dimensional central limit theory for hyper-rectangles. For $l=\mathtt{SSE}$, additional approximation error is controlled using Nazarov’s inequality (Nazarov, 2003).
@@
Step 4: Lower bound on imbalanced split probability. The unconditional Gaussian processes obtained in Step 3 correspond to the squared norm of a univariate ($l\in\{\mathtt{DIM},\mathtt{IPW}\}$) or bivariate ($l=\mathtt{SSE}$) Ornstein–Uhlenbeck process, with a one-to-one transformation between split index for the tree and time for the O-U process (Csörgö and Révész, 1981; Göing-Jaeschke and Yor, 2003). The Darling–Erdös theorem (Eicker, 1979; Horváth, 1993) then characterizes the distribution of the maximum of this process over an interval. Combining this result with the Gaussian correlation inequality (Latała and Matlak, 2017, Remark 3(i)) yields the lower bound in (6), which in turn determines the effective sample sizes of the child nodes.
The remaining arguments apply these insights recursively to deeper trees, allowing us to characterize the concentration properties of the resulting CATE estimators.
@@ 5 $\mathbf{X}$-Adaptivity and Inconsistency
The estimators considered in Theorem 1 either employ the full sample throughout their construction or rely on a two-sample independent split (honesty), where one subsample is used to construct the tree and the other is used to estimate the conditional average treatment effects. As discussed in Devroye et al. (2013) and references therein, $\mathbf{X}$-adaptivity offers a middle ground between these two data-usage designs: the tree construction and the final estimation step share the same covariates but use different outcome variables, so that the two subsamples are independent conditional on the covariates.
We leverage the idea of $\mathbf{X}$-adaptivity and study causal tree estimators in which the outcome variable and treatment indicator are independent across all stages of the tree construction and the final CATE estimation step, while the same covariates are used throughout. This $\mathbf{X}$-adaptive design is of theoretical interest because it provides a bridge between no-sample-splitting and honesty. The following definition formalizes the construction of the corresponding $\mathbf{X}$-adaptive causal tree estimators.
@@
    The dataset $\mathcal{D}$ is divided into $K+1$ datasets $(\mathcal{D}_{\mathsf{T}_{1}},\ldots,\mathcal{D}_{\mathsf{T}_{K}},\mathcal{D}_{\tau})$, with sample sizes given by $(n_{\mathsf{T}_{1}},\ldots,n_{\mathsf{T}_{K}},n_{\tau})$, respectively, and satisfying $n_{\mathsf{T}_{1}}=\cdots=n_{\mathsf{T}_{K}}=n_{\tau}$ (possibly after randomly discarding at most $n\mod K$ observations). For each of the datasets $\mathcal{D}_{j}=\{(y_{i},d_{i},\mathbf{x}_{i}^{\top}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$, $j=1,\ldots,K$, replace $\{(y_{i},d_{i}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$ with independent copies $\{(\tilde{y}_{i},\tilde{d}_{i}):i=1,\ldots,n_{\mathsf{T}_{j}}\}$, while keeping the same $\{\mathbf{x}_{i}:i=1,\ldots,n_{\mathsf{T}_{j}}\}$.
@@
    The maximal decision tree of depth $K$, denoted $\mathsf{T}^{l}_{K}(\mathcal{D}_{\mathsf{T}_{1}},\cdots,\mathcal{D}_{\mathsf{T}_{K}})$, is constructed by iterating $K$ times the splitting procedure $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$ in Definition 2, splitting all terminal nodes until either (i) the node contains a single data point $(y_{i},d_{i},\mathbf{x}_{i}^{\top})$, or (ii) the input values $\mathbf{x}_{i}$ and/or all $(d_{i},y_{i})$ within the node are the same.
@@
As in the previous cases, if the distributions of $\varepsilon_{i}(0)$ and $\varepsilon_{i}(1)$ are symmetric about zero, then the $\mathbf{X}$-adaptive estimators are unbiased: $\mathbb{E}[\hat{\tau}^{\mathtt{X}}_{l}(\mathbf{x};K)]=\tau$, for $l\in\{\mathtt{DIM},\mathtt{IPW},\mathtt{SSE}\}$. The next theorem shows that even under this intermediate sampling design, causal tree estimators remain uniformly inconsistent.
@@
The theorem establishes uniform inconsistency of the $\mathbf{X}$-adaptive causal tree estimator whenever $K_{n}\gtrsim\log\log n$. To put this rate condition in perspective, note that even for extremely large samples, the quantity $\log(\log(n))$ grows very slowly. For example, if $n/K_{n}\approx 1\;\text{billion}$ then $\log\log(10^{9})\approx 3$. Thus, the inconsistency phenomenon appears even for very shallow trees, with only a few recursive splits.
At the same time, the theorem shows that the integrated mean squared error of a uniformly inconsistent $\mathbf{X}$-adaptive causal tree estimator can still decay at the optimal $\sqrt{n}$ rate, up to polylogarithmic factors. As in the previous results, the performance of causal tree estimators can therefore vary dramatically depending on whether the covariate value $\mathbf{x}$ corresponds to an average region of the covariate space or to a worst-case region.